A bird uses a coordinate system of X to the right and Y upwards. The bird has a vector with an X-component of +12.9 ft and a magnitude of 33.1 ft and knows that the vector has a downward component. This problem DOES need to be written out for your portfolio NOTES IMAGES DISCUSS UNITS STATS HELP Part Description Answer Chk History What is the Y component of the vector? (include units with answer) 13.75 pts, 110% # tries: 0 Show Details Format Check 2% try penalty Hints: 0,0 What is the angle of the vector measured counterclockwise (CCW) from the positive x-axis? Answer must be within 0.5 degrees (include units with answer) 13.75 pts, 110% #tries: 0 Show Details Format Check 290 try penalty Hints: 2,0
Answer #1

magnitude of vector = 33.1 ft

X component of vector=12.9 ft

let magnitude of Y component be y

33.12=12.92+y2

y=30.48 ft

given vector has downward y component

therefore, y component of vector = -30.48 ft

angle of vector measured counterclockwise from X axis = 360-tan-1(30.48/12.9)=292.94o